Using gaussian elimination method find all solutions to the following system of linear equations.

2x₂ + 3x₃ + 4x₄ = 1

x₁ - 3x₂ + 4x₃ + 5x₄ = 2

-3x₁ + 10x₂ - 6x₃ -7x₄ = -4


1
Expert's answer
2021-05-19T12:58:05-0400
"\begin{bmatrix}
 0 & 2 & 3 & 4 & \ \ \ 1 \\
 1 & -3 &4 & 5 & \ \ \ 2 \\
-3 & 10 & -6 & -7 & \ \ \ -4 \\
\end{bmatrix}"



Swap rows 1 and 2


"\begin{bmatrix}
 1 & -3 & 4 & 5 & \ \ \ 2 \\
 0 & 2 & 3 & 4 & \ \ \ 1 \\
-3 & 10 & -6 & -7 & \ \ \ -4 \\
\end{bmatrix}"

"R_3=R_3+3R_1"


"\begin{bmatrix}
 1 & -3 & 4 & 5 & \ \ \ 2 \\
 0 & 2 & 3 & 4 & \ \ \ 1 \\
0 & 1 & 6 & 8 & \ \ \ 2 \\
\end{bmatrix}"



"R_2=R_2\/2"


"\begin{bmatrix}
 1 & -3 & 4 & 5 & \ \ \ 2 \\
 0 & 1 & 3/2 & 2 & \ \ \ 1/2 \\
0 & 1 & 6 & 8 & \ \ \ 2 \\
\end{bmatrix}"


"R_1=R_1+3R_2"


"\begin{bmatrix}
 1 & 0 & 17/2 & 11 & \ \ \ 7/ 2 \\
 0 & 1 & 3/2 & 2 & \ \ \ 1/2 \\
0 & 1 & 6 & 8 & \ \ \ 2 \\
\end{bmatrix}"



"R_3=R_3-R_2"

"\begin{bmatrix}
 1 & 0 & 17/2 & 11 & \ \ \ 7/ 2 \\
 0 & 1 & 3/2 & 2 & \ \ \ 1/2 \\
0 & 0 & 9/2 & 6 & \ \ \ 3/2 \\
\end{bmatrix}"


"R_3=2R_3\/9"


"\begin{bmatrix}
 1 & 0 & 17/2 & 11 & \ \ \ 7/ 2 \\
 0 & 1 & 3/2 & 2 & \ \ \ 1/2 \\
0 & 0 & 1 & 4/3 & \ \ \ 1/3 \\
\end{bmatrix}"


"R_1=R_1-17R_3\/2"


"\begin{bmatrix}
 1 & 0 & 0 & -1/3 & \ \ \ 2/ 3 \\
 0 & 1 & 3/2 & 2 & \ \ \ 1/2 \\
0 & 0 & 1 & 4/3 & \ \ \ 1/3 \\
\end{bmatrix}"


"R_2=R_2-3R_3\/2"

"\begin{bmatrix}
 1 & 0 & 0 & -1/3 & \ \ \ 2/ 3 \\
 0 & 1 & 0 & 0 & \ \ \ 0 \\
0 & 0 & 1 & 4/3 & \ \ \ 1/3 \\
\end{bmatrix}"




"\begin{matrix}
 x_1-\dfrac{1}{3}x_4=\dfrac{2}{3} \\ \\
 x_2=0 \\ \\
 x_3+\dfrac{4}{3}x_4=\dfrac{1}{3} \\ \\
\end{matrix}"

Solution set:

"\begin{matrix}
 x_1=\dfrac{2}{3} +\dfrac{1}{3}x_4 \\ \\
 x_2=0 \\ \\
 x_3=\dfrac{1}{3}-\dfrac{4}{3}x_4 \\ \\
x_4,\ free \\ \\
\end{matrix}"